Assume that the BER is 10-4. What is the probability that the 500-bit frame would be...

  1. Assume that the BER is 10-4.
  1. What is the probability that the 500-bit frame would be received without an error (5 points)

If Stop-and-wait flow control is used, what is the probability that the station must transmit 3 frames before a positive acknowledgement is issued? Assuming no acknowledgement lost
Answer #1

a) Probability of single bit error or BER = 10-4

Number if bits in frame = 500

Probability of receiving frame without an error = (1- BER)number of bits in frame

= ( 1- 10-4) 500

As per HomeworkLib policy, sopviso first part, please post rest of them separately